OpenStudy (anonymous):

The formula r=2√5L can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of length L, in feet. How far will a car skid at 80 mph? The car will skid ? feet.

OpenStudy (anonymous):

is L under root

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[r = 2\sqrt{5L}\] or, \[r = 2\sqrt{5}L\]

OpenStudy (anonymous):

first one or second one

OpenStudy (anonymous):

it looks like the first one.

OpenStudy (anonymous):

ok \[80 = 2\sqrt{5L}\] \[\sqrt{5L} = 40\] \[ 5L = 40^2 = 1600\] \[L = 320 miles\] now convert it into feets

OpenStudy (anonymous):

IDK HOW TO DO THAT

OpenStudy (anonymous):

CAN U JUST GIVE ME THE ANSWER

OpenStudy (anonymous):

1 Mile= 5280 Feet 320 Miles x 5280 feet = 1689600 feet

OpenStudy (anonymous):

it said 1689600 is wrong. :(

OpenStudy (ash2326):

r=2 sqrt(5)L 80=2 sqrt(5)L L=17.88854 miles in feet 94451.511

OpenStudy (anonymous):

it said 320 was the correct answer

OpenStudy (ash2326):

i used the second equation so first must be the one,